OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-5, -1) and is perpendicular to y = x + 2.

OpenStudy (anonymous):

Use the opposite reciprocal of original slope for perpendicular equations and plug it into point slope form which is y-y1=m(x-x1) Opposite reciprocal of 1 is -1 So now plug in points and slope y-(-1)=-1(x-(-5)) this equals y+1=-x-5 in normal line format answer would be y= (-x)-6